OpenStudy (anonymous):

Hemal wanted to make dessert for her family. The recipe made 8 servings, but she only wanted to make enough for 5 people. If the recipe for 8 servings called for 2 cups of flour, how much flour would she need for 5 servings?

OpenStudy (lgbasallote):

i believe this is ratio and proportion problem

OpenStudy (lgbasallote):

8 : 2 = 5 : x solve for x

OpenStudy (lgbasallote):

you are familiar with that notation right?

OpenStudy (anonymous):

not really

OpenStudy (lgbasallote):

what's your lesson?

OpenStudy (anonymous):

its a mixture but its fractions

OpenStudy (anonymous):

If you write what was written above as fractions: 8/2 = 5/x You can solve for x. Do you know how to do that?

OpenStudy (lgbasallote):

are you sure your lesson is not ratio and proportion?

OpenStudy (anonymous):

not really im stuck on fractions

OpenStudy (lgbasallote):

i have no idea how to solve this via fractions except for what i and @ten have said..

OpenStudy (anonymous):

Multilpy both sides of the equation by 2. That will remove it from the left side, and you have 5 * 2 on the right side for the numerator. Does that make sense?

OpenStudy (lgbasallote):

@ten i dont think we should be solving this by ratio and proportion if that's not the lesson...

OpenStudy (anonymous):

so is it 8*2 and 5*2 as the fraction

OpenStudy (anonymous):

2 * (8/2) = 8 The reason you multiply by to is to remove it from the denominator.

OpenStudy (anonymous):

oops by 2

OpenStudy (anonymous):

Now you should have: 8 = 10/x

OpenStudy (anonymous):

so i got 5/16

OpenStudy (anonymous):

Did you get to: 8 = 10/x the 16 is not correct. You shouldn't multiply 2 * 8, that 2 just cancelled out the 2 that was in the denominator.

OpenStudy (anonymous):

what am i doing wrong

OpenStudy (anonymous):

2 (8/2) = 2 * (5/x) Equals 8 = 10/x

OpenStudy (anonymous):

i got 4

OpenStudy (anonymous):

The next step would be to multiply both sides by x to get rid of it in the denominator on the right side: 8x = 10

OpenStudy (anonymous):

Then divide both sides by 8, so that x is by itself: x = 10/8 which can be reduced to x = 5/4

OpenStudy (anonymous):

now i got 1.25

OpenStudy (anonymous):

That's the same as 5/4

OpenStudy (anonymous):

right so i got 1 1/4

OpenStudy (anonymous):

That's correct!